OpenStudy (haleyelizabeth2017):

Which term best describes the solution of the situation represented by the system of inequalities? (Assume that x is greater than or equal to 0 and y is greater than or equal to 0.)

OpenStudy (haleyelizabeth2017):

\[x+2y \le 4\]\[x-y \le 1\]\[f(x,y)=3x+2y\]

hero (hero):

This is the same problem type as before

OpenStudy (haleyelizabeth2017):

It is? I graphed it and found three intersection points, but one is (0,-1). Would that count?

hero (hero):

"Which term" Implies options to Choose from.

hero (hero):

There's no way to help you without you listing the choices.

OpenStudy (haleyelizabeth2017):

Oh, sorry....I was finishing the second problem

OpenStudy (haleyelizabeth2017):

Alternate optimal solutions one optimal solution infeasible unbounded

OpenStudy (haleyelizabeth2017):

I think alternate optimal solutions...?

hero (hero):

Why do you think A is the proper choice?

OpenStudy (haleyelizabeth2017):

Bleh...because there are two feasible options? There are those two intersection points that follow the \(x \ge 0\) and \(y\ge0\) thing...

OpenStudy (haleyelizabeth2017):

I'm not entirely positive.

hero (hero):

There's likel yonly one optimal solution for this one. you can test it which solution is optional by inserting the possible points into the equation that Represents the feasible region.

hero (hero):

There Can only be one feasible region for the given set of inequalities

OpenStudy (haleyelizabeth2017):

Ah, thank you for your help. I gotta get going!

hero (hero):

Sorry for the slow response time . It is difficult to use this site on a tablet.

OpenStudy (haleyelizabeth2017):

It's alright :) I can imagine it would be lol

hero (hero):

My Bluetooth keyboard should be arriving soon. Thaf will make it easier I think

OpenStudy (haleyelizabeth2017):

That's good :) Anyhoo, have a good evening!